# A Planar Huygens Antenna Utilizing Crossed Electric and Magnetic Dipoles

### v1

@@ V Conclusion
This paper proposes a superdirective antenna that utilizes crossed electric and magnetic dipoles, forming a Huygens antenna. A magnetic dipole is induced in the slot when it is placed near a printed dipole that is twice the size of the slot. The induced magnetic dipole interacts with the electric dipole mode of the printed dipole; destructive interference occurs in the backward direction, and constructive interference occurs in the forward direction, resulting in a highly directive pattern. The proposed Huygens’ antenna is designed to operate at 4.5 GHz, achieving a directivity of 8.37 dBi with a front-to-back ratio of 12 dB. The radiation efficiency of the antenna is 97%, corresponding to a gain achieved of 8.27 dBi. The high realized gain achieved from a compact planar structure with a high front-to-back ratio makes it suitable for several high-gain applications.

### v2

@@ V Conclusion
This paper demonstrates a superdirective antenna for secure and ultra-low power wireless communication using crossed electric and magnetic dipoles, forming a Huygens antenna. A magnetic dipole based on a half-wavelength slot is placed near a printed dipole that is twice the size of the slot. The induced magnetic dipole interacts with the electric dipole mode of the printed dipole; destructive interference occurs in the backward direction, and constructive interference occurs in the forward direction, resulting in a highly directive pattern. The proposed Huygens’ antenna is designed to operate at 4.5 GHz, achieving a directivity of 8.37 dBi with a front-to-back ratio of 12 dB. The radiation efficiency of the antenna is 97%, corresponding to a gain of 8.27 dBi. The high realized gain achieved from a compact planar structure with a high front-to-back ratio makes it suitable for several high-gain applications.